When Bill was asked how old he was, he replied: "In two years I would have 2 times the age I was 5 years ago. How old is he?
1
Expert's answer
2011-11-15T10:39:53-0500
let's denote x - current age of Bill
hence we can construct the equation
x+2=2(x-5)
x+2=2x-10
x=12